# Formation and disruption of wide binaries in star clusters revealed by N-body simulations

### 02 Apr 2026 v3

@@ III.3 Binary fraction
Figure 4 panel (a) shows the radial binary fraction for all binary types in the example simulation at different times. The models are initialized as unsegregated, and primordial mass segregation is not present. At 0 Myr, $f_{b,r}$ drops toward cluster center. This is owing to the disruption of wide binaries at the dense center at the beginning. Wide binaries continues to dissolve up to 4 half-mass radius at the age of 300 Myr, consistent with the best fitted long-term disruption timescale $t_{2}$ (equation 1) of wide binaries (Table 1). Our simulation results show good agreement with observations of OCs in the solar neighborhood (Pang et al., 2023), which also found a declining binary fraction toward the center. Around 500 Myr, the binary fraction along the radial direction is flat, when mostly hard binaries survive. In the later stages of evolution, binaries move closer toward the central region (Shu et al., 2021) caused by two-body relaxation. Consequently, the radial binary fraction distribution reaches a peak at the center and declines toward the outer part.
We present the binary fraction of all, close, intermediate, and wide binaries as a function of time in Figure 4 panel (b). In our simulations, the close and intermediate binaries are mainly hard binaries. Their fraction changes little over time. The total binary fraction decreases over time, from 17% to 10% within 10 Myr. This decline is mostly due to the disruption of wide binaries, the fraction of which decreases from 8% at 0 Myr to 5% at 2-3 Myr, and to 2% at 10 Myr. Our model resembles ONC in the solar neighborhood, which is observed to contain $\approx 5\%$ wide binaries at the age of 2.5 Myr (Jerabkova et al., 2019). Note that our simulations do not include gas expulsion. The cluster initially does not expand significantly, unlike the one undergoing gas removal, as observed in ONC. The expansion lowers the cluster’s density and favors the survival of wide binaries. This may explain that even though ONC has 3 times lower number of stars than our modeled clusters, it retains a similar fraction of wide binaries as our models. In comparison with other much less-massive and older OCs like Alpha Per (85 Myr, Barrado y Navascués et al., 2004), Pleiades (125 Myr, Stauffer et al., 1998), and Praesepe (790 Myr, Brandt and Huang, 2015) from the study of Deacon and Kraus (2020), the wide binary fraction in our models is $\approx 1\%$ lower than theirs. These clusters may experience much severer gas expulsion and have a lower density. The low-density environment is suitable for the survival of wide binaries, which may become a major contributor to the field population of wide binaries.
@@ IV Conclusion
In this paper, we investigate the dynamical evolution of wide binaries in star clusters using direct NN-body simulations. Although wide (soft) binaries are expected to be rapidly disrupted in dense stellar environments, observations show that they persist in the Galactic field and are moderately observed in OCs. These motivated our study of formation, disruption, and long-term survival of wide binaries.
@@
In this work, we provide a coherent physical framework for the evolution of wide binaries in star clusters. We plan to present a deeper analysis of our simulations in terms of the evolution of compact interacting objects and merger products in the upcoming paper (Bissekenov et al. in prep.).
We wish to express our gratitude to the anonymous referee for providing comments and suggestions that helped to improve the quality of this paper. We thank Jakob Stegmann for the discussions that inspired us to do this work. This research has been funded by the Science Committee of the Ministry of Science and Higher Education, Republic of Kazakhstan (Grant No. AP26102895). X.P. acknowledges the financial support of the National Natural Science Foundation of China through grants 12573036, 12233013 and the China Manned Space Program with grant No. CMS-CSST-2025-A08. PB thanks the support from the special programme of the Polish Academy of Sciences and the U.S. National Academy of Sciences under the Long-term programme to support Ukrainian research teams grant No. PAN.BFB.S.BWZ.329.022.2023. PB also gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant No.  PLG/2026/019243. R.S. acknowledges support from NAOC International Cooperation Office in 2023, 2024, and 2025, from Chinese Academy of Sciences President’s International Fellowship Initiative for Visiting Scientists (PIFI, grant No. 2026PVA0089), from the National Natural Science Foundation of China (NSFC) under grant No. 12473017, and from German Science Foundation (DFG) grant Sp 345/24-1.

### 14 Apr 2026 v4

@@ III.3 Binary fraction
Figure 4 panel (a) shows the radial binary fraction for all binary types in the example simulation at different times. The models are initialized as unsegregated, and primordial mass segregation is not present. At 0 Myr, $f_{b,r}$ drops toward cluster center. This is owing to the disruption of wide binaries at the dense center at the beginning. Wide binaries continue to dissolve up to 4 half-mass radius at the age of 300 Myr, consistent with the best-fitted long-term disruption timescale $t_{2}$ (equation 1) of wide binaries (Table 1). Our simulation results show good agreement with observations of OCs in the solar neighborhood (Pang et al., 2023), which also found a declining binary fraction toward the center. Around 500 Myr, the binary fraction along the radial direction is flat, when mostly hard binaries survive. In the later stages of evolution, binaries move closer toward the central region (Shu et al., 2021) caused by two-body relaxation. Consequently, the radial binary fraction distribution reaches a peak at the center and declines toward the outer part.
We present the binary fraction of all, close, intermediate, and wide binaries as a function of time in Figure 4 panel (b). In our simulations, the close and intermediate binaries are mainly hard binaries. Their fraction changes little over time. The total binary fraction decreases over time, from 17% to 10% within 10 Myr. This decline is mostly due to the disruption of wide binaries, the fraction of which decreases from 8% at 0 Myr to 5% at 2-3 Myr, and to 2% at 10 Myr. Our model resembles ONC in the solar neighborhood, which is observed to contain $\approx 5\%$ wide binaries at the age of 2.5 Myr (Jerabkova et al., 2019). Note that our simulations do not include gas expulsion. The cluster initially does not expand significantly, unlike the one undergoing gas removal, as observed in ONC. The expansion lowers the cluster’s density and favors the survival of wide binaries. This may explain that even though ONC has 3 times lower number of stars than our modeled clusters, it retains a similar fraction of wide binaries as our models. In comparison with other much less massive and older OCs like Alpha Per (85 Myr; Barrado y Navascués et al., 2004), Pleiades (125 Myr; Stauffer et al., 1998), and Praesepe (790 Myr; Brandt and Huang, 2015) from the study of Deacon and Kraus (2020), the wide binary fraction in our models is $\approx 1\%$ lower than theirs. These clusters may experience much more severe gas expulsion and have a lower density. The low-density environment is suitable for the survival of wide binaries, which may become a major contributor to the field population of wide binaries.
@@ IV Conclusion
In this Letter, we investigate the dynamical evolution of wide binaries in star clusters using direct $N$-body simulations. Although wide (soft) binaries are expected to be rapidly disrupted in dense stellar environments, observations show that they persist in the Galactic field and are moderately observed in OCs. These motivated our study of formation, disruption, and long-term survival of wide binaries.
@@
In this work, we provide a coherent physical framework for the evolution of wide binaries in star clusters. We plan to present a deeper analysis of our simulations in terms of the evolution of compact interacting objects and merger products in the upcoming paper (A. Bissekenov et al. 2026, in preparation).
We wish to express our gratitude to the anonymous referee for providing comments and suggestions that helped to improve the quality of this Letter. We thank Jakob Stegmann for the discussions that inspired us to do this work. This research has been funded by the Science Committee of the Ministry of Science and Higher Education, Republic of Kazakhstan (grant No. AP26102895). X.P. acknowledges the financial support of the National Natural Science Foundation of China through grants 12573036 and 12233013, and the China Manned Space Program with grant No. CMS-CSST-2025-A08. P.B. thanks the support from the special programme of the Polish Academy of Sciences and the U.S. National Academy of Sciences under the Long-term programme to support Ukrainian research teams grant No. PAN.BFB.S.BWZ.329.022.2023. P.B. also gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant No.  PLG/2026/019243. R.S. acknowledges support from NAOC International Cooperation Office in 2023, 2024, and 2025, from Chinese Academy of Sciences President’s International Fellowship Initiative for Visiting Scientists (PIFI, grant No. 2026PVA0089), from the National Natural Science Foundation of China (NSFC) under grant No. 12473017, and from German Science Foundation (DFG) grant Sp 345/24-1.